athermia000:

The work of a student trying to solve the equation 2(2x − 3) = 6 + 2x + 4 is shown below: Step 1: 2(2x − 3) = 6 + 2x + 4 Step 2: 4x − 3 = 10 + 2x Step 3: 4x − 2x = 10 + 3 Step 4: 2x = 13 Step 5: x = 6.5 In which step did the student first make an error and what is the correct step? can u guys pls help me

supie:

The equation was given as\[2(2x − 3) = 6 + 2x + 4\] The steps he took were Step 1: \[2(2x − 3) = 6 + 2x + 4\] Step 2: \[4x − 3 = 10 + 2x\] Step 3: \[4x − 2x = 10 + 3\] Step 4: \[2x = 13\] Step 5: \[x = 6.5\] Now let's see the difference in his steps, and my steps. Step 1 : \[2(2x − 3) = 6 + 2x + 4\] Step 2 : \[4 x - 6 = 10 + 2x\] Step 3 :\[ 4x - 2x = 10 + 6\] Step 4 : \[2x = 16\] Step 5 : \[x = 8\] Which step does it look like he messed up on?

athermia000:

witch one

supie:

I'm asking you which one did he make his first mistake on. We cant give you the answer. If you mess up I will guide you and correct you.

athermia000:

ok

athermia000:

fr

athermia000:

will its is a test

athermia000:

ya

athermia000:

oh ok its my firths time here soi didnt know